OpenStudy (laddiusmaximus):

calculate g(x), where g(x) is the inverse of f(x). f(x)=x−3 calculate g(x), where g(x) is the inverse of f(x). f(x)=x−3 @Mathematics

OpenStudy (amistre64):

f(g(x)) = x then

jimthompson5910 (jim_thompson5910):

f(x)=x−3 y=x−3 x=y−3 x+3 = y y = x+3 So the inverse is g(x) = x+3

OpenStudy (amistre64):

g(x) - 3 = x when g(x) = x+3

OpenStudy (anonymous):

@amistre can u help me after ur done \

OpenStudy (laddiusmaximus):

I thought it was too, but when I enter it as answer it says its wrong

jimthompson5910 (jim_thompson5910):

well there must be a typo somewhere

OpenStudy (laddiusmaximus):

it asks me to use theorem 1 from the text to solve it

jimthompson5910 (jim_thompson5910):

and what is theorem 1 in the text?

OpenStudy (laddiusmaximus):

g'(b)= 1/f'(g(b))

OpenStudy (amistre64):

and what does the " ' " ticl mean here?

OpenStudy (laddiusmaximus):

g prime and f prime

OpenStudy (laddiusmaximus):

derivative of.

OpenStudy (amistre64):

i dont understand how that plays into the question really; id have to read your material

OpenStudy (laddiusmaximus):

yeah me neither, hence my confusion.

OpenStudy (laddiusmaximus):

Use Theorem 1 from the text to calculate g(x), where g(x) is the inverse of f(x). f(x)=x−3

OpenStudy (laddiusmaximus):

g'(b)=(1)/f'(g(b))

OpenStudy (amistre64):

[f(g(x))]' = f'(g(x)) g'(x) f'(x) = 1; g(x) = 1 so g'(x) = 0 .... 1/(f'(1)) ..... jibberjabber is all i see

OpenStudy (amistre64):

\[\frac{dy}{dx}=\frac{1}{dx/dy}\] is a thrm

OpenStudy (amistre64):

f'(g(x)) g'(x) = 1 from the chain rule g'(x) = 1/f'(g(x))

OpenStudy (amistre64):

is where its from

OpenStudy (laddiusmaximus):

\[g'(b)=1\div f'(g(b))\]

OpenStudy (amistre64):

f(x) = x-3 ; g(x) = x+3 f'(x) = 1 ; g'(x) = 1 g'(x) = 1/f'(g(x)) 1 = 1/f'(x+3) f'(x+3) = 1 its gotta be something to do with these

OpenStudy (amistre64):

joemath might remember it

OpenStudy (laddiusmaximus):

yeah the text really confused me.